Calculate the chloride ion molarity if 5.382 g of AlCl3 is dissolved in 250.00 mL of water
Answer #1

Cl-

m = 5.382 g of AlCl3

calculate moles of AlCl first

MW of Alcl3 = 133.34

mol = mass/MW = 5.382/133.34 = 0.0403629

1 mol of AlCl3 contains 3 mol of Cl, therefore

3*0.0403629 = 0.1210887 mol of Cl- in solution

mut we require molarity

M = mol/L

then

M = 0.1210887 / (0.25) = 0.4843548 M of Cl- ions